Helen invested 40,000 and received a total of 2300.00 in interest after one year. If part of the money returned 5% and the remainder 8% then how much did she invest at each rate?
Answer by macston(5194)

.
F=amount invested at 5%; E=amount invested at 8%
.
F+E=40000
E=40000-F Use this to substitute for E below.
.
0.05F+0.08E=2300 Substitute for E.
0.05F+0.08(40000-F)=2300
0.05F+3200-0.08F=2300 Subtract 3200 from each side.
-0.03F=-900 Divide each side by -0.03.
F=30000 ANSWER 1: She invested 30000 at 5%.
.
E=40000-F
E=40000-30000
E=10000 ANSWER 2: She invested 10000 at 8%.
.
CHECK:
0.05F+0.08E=2300
0.05(30000)+0.08(10000)=2300
1500+800=23==
2300=2300